OpenStudy (anonymous):

in the formula N=le^kt, N is the number of items in term of an initial population I at a given time t and k is a growth constant equal to the percent of growth per unit time. there are currently 61 million car in a certain country increasing by 3.9% annually. how many years will it take for this country to have 82 million cars round to nearest year. plz help meeee

OpenStudy (ranga):

Plug in the numbers and solve for t. N = 82 * 10^6 I = 61 * 10^6 k = 3.9

OpenStudy (anonymous):

plug where

OpenStudy (anonymous):

oh i got it thanks you ranga sir

OpenStudy (ranga):

you are welcome.

OpenStudy (ranga):

Did you get the correct answer?

OpenStudy (anonymous):

i already did that i posted wrong question but thanks you sir

OpenStudy (ranga):

Okay, but I think k should be 0.039 in the formula because I think in the problem they incorrectly state that: "k is a growth constant equal to the percent of growth per unit time" It should be percent of growth in decimal. So 3.9% = 0.039 in decimal.

OpenStudy (anonymous):

oh k thanks you